OpenStudy (anonymous):

I am stuck on this word problem and need help! The official population of Baton Rogue, Louisiana, was 219,531 in 1990 and 227,818 in 2000. a) What was the average rate of change in the pop. from 1990 to 2000? b) Write a linear equation that models the pop. (P) of Baton Rogue from 1990 to 2000. Let t represent the number of years since 1990.

OpenStudy (wolf1728):

I'm looking at this right now.

OpenStudy (anonymous):

is this the problem you need help on?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright do you want the answers or the explanation?

OpenStudy (anonymous):

Both if you're willing

OpenStudy (anonymous):

I tried figuring this out but couldnt so asked my sister for help a. 834 people per year b. P = 834t + 219,478 these are the answers i got from asking my sister

OpenStudy (anonymous):

Okay. That looks like it's right whereas I couldn't get anything. thank you

OpenStudy (wolf1728):

I'm working on it as if the change is logarithmic as opposed to linear.

OpenStudy (anonymous):

Okay, thank you

OpenStudy (wolf1728):

I noticed you left for a bit - do you want to stick around?

OpenStudy (anonymous):

Yes. Just trying to figure out this brain rattler

OpenStudy (anonymous):

There is also another question, but I do not know how to ask it. I just started this website..

OpenStudy (wolf1728):

The population changes by a factor of 1.00373 every year. So population in 1991 = 1990 population * (1.00373) = 220,350 1992 population = 219,531 * (1.00373)^2 = 221,172

OpenStudy (anonymous):

The answer that turquoisetuna gave me is similar to what I was thinking.. I just need to write an equation.

OpenStudy (wolf1728):

Hmm, I also notice that the problem states that you are to "write a linear equation". My solution involves a logarithmic progression (which is also more complex). Maybe you should stick with what turqoisetuna said.

OpenStudy (anonymous):

Okay, well thank you for trying to help. It is very much appreciated

OpenStudy (wolf1728):

I can work on the equation for you. That should be easy.

OpenStudy (anonymous):

I have always been terrible at word problems. So thank you both for helping. There is also a third part to this question though that says to estimate what the pop. will be in 2010 using an equation you come up with...

OpenStudy (wolf1728):

227,818 -219,531 = 8,287 which is the population change for 10 years So population changes 829 people every year. The equation is population = 219,531 + (829*t) population in 1995 = 219,531 + (829*5) population in 1995 = 219,531 + 4,145 population in 1995 = 223,676

OpenStudy (anonymous):

Thank you! (tuna too)